OpenStudy (anonymous):

What is the slope of the line that passes through the pair of points (2, 9), and (5, 12)? A. 0 B. 1 C. 3 D. undefined

OpenStudy (anonymous):

Consider the slope formula m = (y2 - y1) / (x2 - x1). Give it a shot.

OpenStudy (anonymous):

0?

OpenStudy (anonymous):

Let's let (x1, y1) = (2, 9) and (x2, y2) = (5, 12). (y2 - y1) / (x2 - x1) = (12 - 9) / (5 - 2) = 3/3 = 1 = slope.

OpenStudy (anonymous):

Aah! I see why i got it wrong now. Thank you!